# Careful finite element simulations of cold rolling with accurate through-thickness resolution and prediction of residual stress

### v1

@@ 5 Comparison to analytical models
In this section, we compare FE simulations to the rigid-perfectly-plastic analytical model of Erfanian et al. [13]. We also compute the residual stress in the sheet after rolling and investigate its connection to the oscillatory pattern in stress and strain quantities in the roll gap. This connection is likely to have a strong relationship with curvature in asymmetric rolling, a phenomenon that is often incorrectly predicted [63].
5.1 Comparing FE outputs to the analytical results
Figure 7 shows shear stress contour plots ($\mathrm{M}\mathrm{P}\mathrm{a}$) comparing the FE outputs with the analytical results of Erfanian et al. [13]. Three roll-gap aspect ratios are investigated: $\varepsilon=0.175$, $\varepsilon=0.125$ and $\varepsilon=0.075$. As Erfanian et al. [13] do not consider deformations outside the roll gap, our attention here is restricted to values of $x/L$ between 0 and 1 (see Figure 1 for domain definition). For each value of $\varepsilon$, an oscillatory pattern is predicted by both the FE simulations and the analytical model. The number and position of the predicted shear stress lobes are in agreement for both methods. Similar to Minton [43], the number of lobes increases with decreasing $\varepsilon$. The magnitude of shear stress at various through-thickness positions is also noted to be, in general, similar. There are however two notable differences between the two approaches. Firstly, the lobes in the FE results have a more rounded profile compared to the analytical outputs, particularly for larger values of $\varepsilon$. As $\varepsilon$ decreases the lobe profiles become more similar. This is unsurprising as the analytical model of Erfanian et al. [13] is only valid in the limit $\varepsilon\rightarrow 0$.
@@ 5.1 Comparing FE outputs to the analytical results
Figure 7: Contour plots of shear stress results from simulations with $N_{e}=30$ elements. The left panel (a-c) shows FE simulations and the right panel (d-f) shows results from the analytical model of Erfanian et al. [13], for three different roll-gap aspect ratios: $\varepsilon=0.175$ (top row, $R=130.6\,$\mathrm{m}\mathrm{m}$$), $\varepsilon=0.125$ (middle row, $R=257.45\,$\mathrm{m}\mathrm{m}$$) and $\varepsilon=0.075$ (bottom row, $R=710\,$\mathrm{m}\mathrm{m}$$). Data is presented for the roll gap only, between $x/L=0$ and $x/L=1$, in all cases since Erfanian et al. [13] do not consider deformations outside the roll gap.
@@
The second difference between the two approaches is the predicted shear profile between the neutral point and the end of the roll gap (i.e., $x_{N}<x/L<1$). The neutral point $x_{N}$ is defined as the position at which friction changes direction on the surface. In Figure 7 the neutral point is observed at $x/L\approx 0.6913$ in all contour plots. In the FE results, shear stress on the surface varies in a smooth fashion while the analytical model demonstrates a distinct discontinuity in shear stress. The discontinuous change in shear stress seen in the analytical model is due to the slipping Coulomb friction law that is employed. In reality, material sticks to the roller around the neutral point and shear stress changes sign smoothly [23, 14]. This sticking region around the neutral point is predicted in our FE results but not in the rigid-plastic analytical model.
@@
Figure 8: Shear stress along the top surface of the deformed sheet from the analytical model by Erfanian et al. [13] and from simulations with different Young’s modulus ($E$) and elastic slip tolerance ($\gamma$) values. The default values are $E=206.3\,$\mathrm{G}\mathrm{P}\mathrm{a}$$ and $\gamma=0.005$. All simulations have radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
@@
Figure 9 shows von Mises stress and PEEQ rate contour plots, along with the sheet’s surface and central velocity magnitudes, from the two simulations with Young’s moduli $E=206.3\,$\mathrm{G}\mathrm{P}\mathrm{a}$$ and $E=1237.8\,$\mathrm{G}\mathrm{P}\mathrm{a}$$ ($6E$). The PEEQ rate results were calculated in a similar manner to the velocities (see Section 2 of the supplementary information provided). It is interesting to note the clearly non-vertical boundaries between the sub-yield regions and the yielding regions around the roll gap entrance and exit in Figures 9(a) and 9(d). Many analytical models make the assumption of vertical boundaries between these regions. Another interesting point to note is that the end of the plastic deformation zone is closer to $x/L=1$ in Figure 9(d) than in Figure 9(a), indicating less elastic springback for a higher Young’s modulus. This is unsurprising since elastic strains decrease as $E\rightarrow\infty$.
There are differences between Figures 9(b) and 9(e), the two PEEQ rate contour plots, but the general trends are the same. A clear oscillatory pattern exists in the PEEQ rate plots, and regions of zero-valued PEEQ rate periodically appear at the sheet’s surface and centre. Regions of zero PEEQ rate indicate that no plastic flow is occurring in these regions. However, the von Mises results do not indicate that the stress is significantly below the yielding value in the neighbourhood of any of these no-plastic-flow zones. Therefore we refer to these regions as no-plastic-flow regions as opposed to “sub-yield” regions. The surface and central velocity plots in Figures 9(c) and 9(f) also show clear oscillatory behaviours. Each region of constant velocity at the sheet’s surface or centre coincides with a no-plastic-flow region at the corresponding sheet location. This is plausible since constant velocity implies a rigid-body motion. We also see a sticking region in both surface velocity outputs around the neutral point, where the speed of the surface of the sheet is equal to the roller’s tangential speed. This allows the shear stress to smoothly change from positive to negative on the surface as the sheet sticks to the rollers for a small portion of the roll gap. A smaller sticking region is observed for the $6E$ simulation. We hypothesise that this sticking region would continue to shrink, and hence tend toward the analytical output, in the limit as $E\rightarrow\infty$, but given convergence issues for higher $E$ values than are given here, this remains a hypothesis.
We have shown that by decreasing the value of $\gamma$ and increasing the value of $E$, the FE results become more similar to the analytical results of Erfanian et al. [13]. Therefore, we hypothesise that by taking the limit as $\gamma\rightarrow 0$ and as $E\rightarrow\infty$, it would be possible to achieve much better agreement between the FE results and the rigid-plastic shear stress results given by Erfanian et al. [13] around the neutral point. However, decreasing the slip tolerance $\gamma$ and increasing the Young’s modulus $E$ are computationally intensive tasks that suffer from convergence issues when more extreme values are attempted. The limitations of the analytical model of Erfanian et al. [13] in terms of accurately predicting the shear stress distribution are therefore clear: (1) The shear lobe shape deviates from FE predictions as the roll-gap aspect ratio $\varepsilon$ increases from zero and (2) the material and friction models are inadequate for capturing sticking behaviours and the smoothly-varying shear stress around the neutral point.
Figure 9: Comparison of FE simulations for two Young’s moduli. The left panel (a–c) shows results from the FE simulation with the original Young’s modulus ($E=206.3\,$\mathrm{G}\mathrm{P}\mathrm{a}$$). The right panel (d–f) shows results for a Young’s modulus six times greater than that of (a–c). Panels (a) and (d) show the von Mises contour plots for each simulation, while panels (b) and (e) show the PEEQ rate contour plots. Panels (c) and (f) show the sheet’s surface and central velocity magnitudes (solid and dashed lines respectively) for each simulation, and the roller’s tangential velocity (horizontal black dotted line). The PEEQ rate and velocity values are obtained through postprocessing of the PEEQ values and displacements respectively (see Section 2 of the supplementary material). Both simulations employ $N_{e}=30$ CPE4R elements, a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
5.2 Residual stress
Here we display the residual stress in the sheet after rolling and investigate its connection to the through-thickness variation of the shear stress in the roll gap. To highlight this connection, we consider a symmetric simulation with a roller radius $R=32\,$\mathrm{m}\mathrm{m}$$, which leads to a roll-gap aspect ratio of $\varepsilon=0.354$, and all other parameters are the same as described in Section 3. We measure the von Mises stress and the quantities $p\pm 2\theta Y/\sqrt{3}$, where $p=-0.5\left(\sigma_{xx}+\sigma_{zz}\right)$ is the hydrostatic pressure and $\theta$ is the anticlockwise rotation angle of the stress element from the positive $x$-axis for which the stress is in a state of pure shear. The von Mises stress in the sheet after it passes through the roll gap can be used to give an indication of residual stress, while the quantities $p\pm 2\theta Y/\sqrt{3}$ are of interest because these quantities are constant along two orthogonal sets of characteristic lines for plastic deformation, and the direction of these characteristic lines coincide with the lines of maximum shear stress (slip lines) when plastic deformation is occurring (see Section 5 of the supplementary material for more details).
Figure 10 shows a contour plot of von Mises stress and of $p\pm 2\theta Y/\sqrt{3}$ ($\mathrm{M}\mathrm{P}\mathrm{a}$) from a symmetric simulation with $N_{e}=30$. The right-end $x/L$ value in the plot is large and the von Mises stress towards the right end of the contour plot can therefore give an indication of residual stress since it is significantly far from the external loading (i.e., the rollers). The distinct centre and outer von Mises zones in the sheet after the roll gap in Figure 10(a) correspond to zones of lengthwise compression and tension respectively. This is consistent with longitudinal stress results reported by Tadić et al. [45]. The length–height ratio of these contour plots is set to the true ratio used in the simulation. This allows for interpretation of the direction of the contour lines in Figure 10(b). For large portions of the roll gap, these contour lines meet each other at $90^{\circ}$, indicating that plastic deformation is occurring. The sub-yield are easily identified by comparing against the von Mises contour plot in Figure 10(a). Recall the initial yield stress is $Y=477.2\,$\mathrm{M}\mathrm{P}\mathrm{a}$$, and thus the regions of the sheet with von Mises values significantly less than this value are sub-yield regions. In these sub-yield regions the $p\pm 2\theta Y/\sqrt{3}$ contour lines are not orthogonal, and thus cannot be called slip lines in these regions. The contour lines are only slip lines in regions of plastic deformation. Another interesting observation from the von Mises contour plot in Figure 10(a) is that there is a significant amount of residual stress under the top surface (and above the bottom surface) of the sheet for $x/L>1$. The $p\pm 2\theta Y/\sqrt{3}$ contour lines oscillate through the roll gap in a similar manner to the shear lobes discussed earlier, which is unsurprising given they line up with the directions of maximum shear stress during plastic flow. After oscillating through the roll gap, the $p\pm 2\theta Y/\sqrt{3}$ contour lines also accumulate within a thin layer that roughly lines up with these regions of residual von Mises stress. This indicates that a strong relationship exists between the oscillations in shear stress in the roll gap and residual stress outside the roll gap.
Figure 10: Contour plots of (a) von Mises and (b) $p\pm 2\theta Y/\sqrt{3}$, all given in $\mathrm{M}\mathrm{P}\mathrm{a}$, from a symmetric implicit simulation with $N_{e}=30$ CPE4R elements. The radius is $R=32\,$\mathrm{m}\mathrm{m}$$ which gives a roll-gap aspect ratio of $\varepsilon=0.354$. The tangential velocity of the rollers is $R\Omega=1287.25\,$\mathrm{m}\mathrm{m}\,\mathrm{s}^{-1}$$.
5.3 Asymmetric rolling
We propose that the connection between shear stress in the roll gap and residual stress significantly far away from the roll gap can be used for curvature prediction in asymmetric rolling. For this investigation, we consider an asymmetric simulation with a roller radius $R=32\,$\mathrm{m}\mathrm{m}$$, where the top roller rotates 2.5% faster than the bottom roller. The bottom roller rotates with the same tangential velocity as before ($R\Omega=1287.25\,$\mathrm{m}\mathrm{m}\,\mathrm{s}^{-1}$$), and all other parameters are the same as described in Section 3.
The asymmetric results shown in Figure 11 are obtained via a simulation with $402N_{e}=$ 40. We use this mesh density here as the equivalent of the mesh size used in previous simulations ($602N_{e}=$ 60) was unachievable due to the whole sheet having to be discretised in the asymmetric simulation as opposed to just the top half. Large amounts of residual stress are observed in the asymmetric von Mises results in Figure 11(a), similar to what was noted in the symmetric case (Figure 10(a)). Once again the contour lines of $p\pm 2\theta Y/\sqrt{3}$ accumulate within a thin layer that lines up with these regions of residual von Mises stress. The magnitude of the residual stress in the sheet for $x/L>1$ is greater towards the bottom of the sheet, and the sheet’s curvature is towards the slower roller for this specific set of parameters. We do not suggest that curvature is always towards the slower roller, because as pointed out by Minton and Brambley [63], many authors have given contradicting prediction rules for curvature during asymmetric rolling. Rather we hypothesise that the oscillatory pattern in stress and strain quantities presented here, which is largely absent in the literature, has a strong influence on the accurate prediction of the residual stress for $x/L>1$, which may have a significant impact on predicting curvature in asymmetric rolling. This is an important aspect of modelling rolling since unintentional asymmetries cause turn-up and turn-down in sheet rolling which can halt the process or even damage the rolling table and mills and since asymmetries are intentionally exploited in asymmetrical rolling to improve process efficiency and produce curved products [63].
Figure 11: Contour plots of (a) von Mises and (b) $p\pm 2\theta Y/\sqrt{3}$, given in $\mathrm{M}\mathrm{P}\mathrm{a}$, from an implicit simulation with $402N_{e}=$ 40 CPE4R elements and asymmetric roller speeds. The radius is $R=32\,$\mathrm{m}\mathrm{m}$$ which gives a roll-gap aspect ratio of $\varepsilon=0.354$. The top roller rotates 2.5% faster than the bottom roller. The tangential velocity of the bottom roller is $R\Omega=1287.25\,$\mathrm{m}\mathrm{m}\,\mathrm{s}^{-1}$$.

### v2

@@ 5 Numerical results
In this section, we present some notable outputs from our FE simulations, highlighting the emergence of a deterministic oscillatory pattern in the shear, PEEQ rate and velocity profiles of metal sheets during cold rolling. We also compute the residual stress in the sheet after rolling and investigate its connection to the oscillatory pattern in stress and strain quantities in the roll gap. This connection is likely to have a strong relationship with curvature in asymmetric rolling, a phenomenon that is often incorrectly predicted [70].
5.1 Oscillations
5.1.1 Shear lobes
@@
Figure 9: Contour plots of shear stress results from simulations with $N_{e}=30$ elements, where the symmetry axis is located at $z/h_{0}=0$ (see Figure 1). The left panel (a-c) shows FE simulations and the right panel (d-f) shows results from the analytical model of Erfanian et al. [21], for three different $\varepsilon$ values: $\varepsilon=0.175$ (top row, $R=130.6\,$\mathrm{m}\mathrm{m}$$), $\varepsilon=0.125$ (middle row, $R=257.45\,$\mathrm{m}\mathrm{m}$$) and $\varepsilon=0.075$ (bottom row, $R=710\,$\mathrm{m}\mathrm{m}$$). Data is presented for the roll gap only, between $x/L=0$ and $x/L=1$, in all cases since Erfanian et al. [21] do not consider deformations outside the roll gap. The neutral point is the horizontal $x/L$ position at which the surface shear stress changes sign.
The left column of Figure 9 shows our FE prediction of shear stress in sheets undergoing rolling in three different roll gaps corresponding to aspect ratios $0.1751/\varepsilon=1/$ 0.175, $1/0.125$ and $1/0.075$. In each of the three cases, we observe a clear oscillatory pattern, with the shear stress varying from $-150$\mathrm{M}\mathrm{P}\mathrm{a}$$ to $+150$\mathrm{M}\mathrm{P}\mathrm{a}$$, and with lobes of positive and negative shear stress appearing with regular spacing on both sides of the sheet mid-surface. The number of lobes increases as $1/\epsilon$ increases, which is consistent with preliminary findings from [18]. These lobes are not typically reported in FE modelling of cold rolling (c.f. Table 1) and could be mistaken for numerical error, but they have also been predicted by recent analytical work (c.f. the right column of Figure 9 and Section 6).
5.1.2 Deforming and non-deforming zones
Figure 10: (a) Von Mises stress, (b) PEEQ rate and (c) velocity profiles for the simulation presented in Figure 9(b) ($R=257.45\,$\mathrm{m}\mathrm{m}$$, $\varepsilon=0.125$). In the contour plots, the symmetry axis is located at $z/h_{0}=0$ (see Figure 1). The PEEQ rate and velocity values are obtained through postprocessing of the PEEQ values and displacements respectively (see B).
Figure 10 shows von Mises stress and PEEQ rate contour plots, along with the sheet’s surface and central velocity magnitudes, for the simulation presented in Figure 9(b) ($R=257.45\,$\mathrm{m}\mathrm{m}$$, $\varepsilon=0.125$). The PEEQ rate results were calculated in a similar manner to the velocities (see B). From Figure 10(a), we see that the von Mises stress does not drop below yield in the interior of the roll gap, as expected. However, the PEEQ rate plot in Figure 10(b) reveals a far richer and more complex layer of dynamics. A clear wave-like structure can be seen, mirroring the oscillatory pattern in shear in Figure 9. The plastic deformation is highly localized in very thin bands, oriented diagonally and oscillating back and forth across the sheet. In the regions between these bands, where the PEEQ rate is zero, the material appears to translate as a solid body, with no plastic flow, even though it is at yield. This is unexpected and may have major implications for hardening and the development of microstructure in the rolled sheet.
We investigate the emergence of rapidly-deforming and non-deforming zones further by plotting the velocity magnitude at the surface and centre of the sheet in Figure 10(c) (solid and dashed curves, respectively). The speed of the roller (black dotted line) is also shown for comparison. According to these curves, the sheet surface initially accelerates when it enters the roll gap and comes into contact with the roller, as expected from Coulomb friction, while the sheet centreline remains at its initial speed. As shown in Figure 10(c) the velocity at the sheet centre does not increase until $x/L\approx 0.1$ while the sheet surface velocity begins to increase at $x/L\approx 0$. From Figure 9 we know that the shear stress has a non-linear distribution through the sheet thickness, and this delayed response to velocity changes is therefore expected. The velocity of the material at the sheet’s horizontal centre does not change until the information from the surface reaches the sheet’s centre. The acceleration is transmitted to the centre of the sheet via the shear, as shown in Figure 9, so it begins to accelerate later. However, the surface then stops accelerating, in line with the translation-only (i.e., zero PEEQ rate) region observed in the PEEQ rate results. This means that the centreline temporarily accelerates to a faster velocity than the surface. This behaviour is, again, somewhat unexpected, since one might intuitively expect the material velocity to be highest near the sheet surface, where the forcing is imposed. The pattern then repeats, and both regions move through consecutive intervals of acceleration and rigid translation, with the respective velocities leap-frogging each other repeatedly. Furthermore, we observe that one of the velocity plateaus on the surface coincides with the roll velocity, so a “stick” region emerges naturally around the neutral point111 The neutral point is defined as the position at which friction changes direction on the surface (see Figure 9). . This allows the shear stress to smoothly change from positive to negative on the surface as the sheet sticks to the rollers for a small portion of the roll gap, as was observed at the sheet’s surface near the neutral point in the FE results in Figures 9(a)9(c).
5.2 Residual stress
Figure 11: Contour plots of (a) von Mises and (b) $p\pm 2\theta Y/\sqrt{3}$, all given in $\mathrm{M}\mathrm{P}\mathrm{a}$, from a symmetric implicit simulation with $N_{e}=30$ CPE4R elements, where the symmetry axis is located at $z/h_{0}=0$ (see Figure 1). The radius is $R=32\,$\mathrm{m}\mathrm{m}$$ which gives $\varepsilon=0.354$. The tangential velocity of the rollers is $R\Omega=1287.25\,$\mathrm{m}\mathrm{m}\,\mathrm{s}^{-1}$$.
The through-thickness variation of the stress inside the roll gap illustrated in Figure 10 has a knock-on effect on residual stress after rolling. In Figure 11 we consider a symmetric simulation with a roller radius $R=32\,$\mathrm{m}\mathrm{m}$$, which leads to $\varepsilon=0.354$, and all other parameters are the same as described in Section 3. The length–height ratio of these contour plots is set to the true ratio used in the simulation, for ease of interpretation. We measure the von Mises stress and the quantities $p\pm 2\theta Y/\sqrt{3}$, where $p=-0.5\left(\sigma_{xx}+\sigma_{zz}\right)$ is the hydrostatic pressure and $\theta$ is the anticlockwise rotation angle of the stress element from the positive $x$-axis for which the stress is in a state of pure shear. The von Mises stress in the sheet after it passes through the roll gap can be used to give an indication of residual stress. For this reason the plotted domain is extended beyond the roll gap ($x/L>1$). The quantities $p\pm 2\theta Y/\sqrt{3}$ are of interest because these quantities are constant along two orthogonal sets of characteristic lines for plastic deformation, and the direction of these characteristic lines coincide with the lines of maximum shear stress (slip lines), but only when plastic deformation is occurring (see E for more information).
Here, we are concerned with residual stress, seen for $x/L>1.5$ in Figure 11(a), where the material is far from the external loading (i.e., the rollers). We observe distinct zones in the centre and at the edges; cross-checking against the individual stress components reveals that the central zone experiences lengthwise compression while the outer zones are under longitudinal tension. This is consistent with longitudinal stress results reported by Tadić et al. [39]. The residual von Mises stress is highest in narrow horizontal bands in between the sheet surface and the sheet centre. In the portions of the sheet that are at yield in Figure 11(a), the contour lines of $p\pm 2\theta Y/\sqrt{3}$ in Figure 11(b) meet at $90^{\circ}$, as expected theoretically [36]. These are the slip lines. In the sub-yield zone, the contour lines are not slip lines and do not meet at $90^{\circ}$. The $p\pm 2\theta Y/\sqrt{3}$ contour lines oscillate through the roll gap in a similar manner to the shear lobes discussed earlier, which is unsurprising given they line up with the directions of maximum shear stress during plastic flow. After oscillating through the roll gap, the $p\pm 2\theta Y/\sqrt{3}$ contour lines also accumulate within a thin layer that roughly lines up with the regions of largest residual von Mises stress noted above. This indicates that a strong relationship exists between the oscillations in shear stress in the roll gap and residual stress outside the roll gap.
5.3 Asymmetric rolling
We hypothesise that the connection between shear stress in the roll gap and residual stress significantly far away from the roll gap has implications for curvature prediction in asymmetric rolling. Here, by way of a preliminary investigation, we consider an asymmetric simulation with rollers of radius $R=32\,$\mathrm{m}\mathrm{m}$$, where the top roller rotates 2.5% faster than the bottom roller. The bottom roller rotates with the same tangential velocity as before ($R\Omega=1287.25\,$\mathrm{m}\mathrm{m}\,\mathrm{s}^{-1}$$), and all other parameters are the same as described in Section 3.
Figure 12: Contour plots of (a) von Mises and (b) $p\pm 2\theta Y/\sqrt{3}$, given in $\mathrm{M}\mathrm{P}\mathrm{a}$, from an implicit simulation with $402N_{e}=$ 40 CPE4R elements and asymmetric roller speeds. The radius is $R=32\,$\mathrm{m}\mathrm{m}$$ which gives $\varepsilon=0.354$. The top roller rotates 2.5% faster than the bottom roller. The tangential velocity of the bottom roller is $R\Omega=1287.25\,$\mathrm{m}\mathrm{m}\,\mathrm{s}^{-1}$$.
The asymmetric results shown in Figure 12 are obtained via a simulation using $40$ through-thickness elements [^2] The mesh density comparable to the symmetric case above would have required 60 elements through thickness, which was too computationally expensive to be achievable for this asymmetric case. . Large amounts of residual stress are observed in the asymmetric von Mises results in Figure 12(a), similar to what was noted in the symmetric case (Figure 11(a)). Once again the contour lines of $p\pm 2\theta Y/\sqrt{3}$ accumulate within a thin layer that lines up with these regions of largest residual stress. The magnitude of the residual stress in the sheet for $x/L>1$ is greater towards the bottom of the sheet, and the sheet’s curvature is towards the slower roller for this specific set of parameters. We do not suggest that curvature is always towards the slower roller because, as pointed out by Minton and Brambley [70], many authors have given contradicting prediction rules for curvature during asymmetric rolling. Rather we hypothesise that the oscillatory pattern in stress and strain quantities presented here, which is largely absent in the literature, has a strong influence on the accurate prediction of the residual stress for $x/L>1$, which may have a significant impact on predicting curvature in asymmetric rolling. Indeed, the shifting of the oscillatory pattern as parameters are varied likely explains the inconsistent and poorly-understood trends in curvature reported in the literature [70]. This is an important aspect of modelling rolling since unintentional asymmetries cause turn-up and turn-down in sheet rolling which can halt the process or even damage the rolling table and mills and since asymmetries are intentionally exploited in asymmetrical rolling to improve process efficiency and produce curved products [70].
6 Further validation
Many of the numerical observations presented in Section 5 are unexpected or unintuitive, so we provide further validation in this section by comparing with a new analytical method developed using the method of multiple scales by Erfanian et al. [21]. This new model yields a wave-like solution for the stresses in the sheet, where information travels through the sheet along diagonal slip lines and is reflected at the sheet surface, yielding a characteristic zig-zag pattern as illustrated in Figures 9(d)9(f). This multiple-scales model is valid in the limit where the roll gap is very long and thin, i.e. for $\varepsilon\to 0$. We compare our FE predictions of shear stress with the multiple-scales prediction by comparing the left and right columns of Figure 9. As Erfanian et al. [21] do not consider deformations outside the roll gap, our attention here is restricted to values of $x/L$ between 0 and 1 (see Figure 1 for the domain definition). Both methods predict repeating lobes of high and low shear stress, as outlined above. The number and position of lobes predicted by the two methods are in agreement for each value of $\varepsilon$, and the magnitude of shear stress at various through-thickness positions is, in general, comparable. There are some differences between the predictions from the two methods, which we outline and explain now.
6.1 Through-thickness shear profile
@@
One notable difference between the FE and analytic models is in the shear profiles, which can be seen between the neutral point and the end of the roll gap. In Figure 9 the neutral point is observed in the range $x\approx 0.6L\text{--}0.75L$ in all contour plots as the $x$-position at which the shear changes sign on the surface of the sheet. In the FE results, shear stress on the surface varies in a smooth fashion while the analytical model demonstrates a distinct discontinuity in shear stress. The discontinuous change in shear stress seen in the analytical model is due to the slipping Coulomb friction law that is employed. In reality, material sticks to the roller around the neutral point (as shown in Figure 10(c) above) and shear stress changes sign smoothly [31, 30]. This sticking region around the neutral point is predicted in our FE results, but not in the rigid-plastic analytical model.
6.2 Stick–slip region
@@
Figure 13: Shear stress along the top surface of the deformed sheet from the analytical model by Erfanian et al. [21] and from simulations with different Young’s modulus ($E$) and elastic slip tolerance ($\gamma$) values. The default values are $E=206.3\,$\mathrm{G}\mathrm{P}\mathrm{a}$$ and $\gamma=0.005$. All simulations have radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and $\varepsilon=0.125$.